# Panoramic Out-of-Distribution Segmentation for Autonomous Driving

### v2

@@ III Methodology
In our work, to promote a more comprehensive omnidirectional perception and leverage the advantages of panoramic broader-FoV, a novel panoramic perception task, PanOoS, is introduced. In response to the challenges posed by complex backgrounds and pixel distortions, we analyze the model architecture and fine-tuning strategies, ultimately proposing the first mask-based out-of-distribution segmentation model specifically designed for the established PanOoS task.
@@ III-A Problem Setting
Out-of-distribution segmentation and anomaly segmentation are inherently similar; however, more specifically, out-of-distribution segmentation encompasses anomaly segmentation. The primary distinction lies in whether the methods can segment each in-distribution category while also identifying outliers. Out-of-distribution segmentation can be achieved in per-pixel semantic segmentation architectures [29, 27, 28, 26, 72] by computing anomaly probability on top of the per-pixel classifier. For the pixel-level class scores $S(x){\in}\mathbb{R}^{K\times H\times W}$ obtained by segmenting the input image $x{\in}\mathbb{R}^{3\times H\times W}$ using a pixel-wise segmentation architecture, the anomaly scores $A(x){\in}\mathbb{R}^{H\times W}$ are computed as follows:
$A(x)=f(S(x)),$ (1)
where $K$ denotes the number of classes for closed-set segmentation, $H$ and $W$ correspond to the height and width of the image, respectively.
Figure 4: Overview of POS. The proposed POS consists of four primary components: an image encoder, a text encoder, a pixel decoder, and a transformer decoder with Prompt-based Restoration Attention (PRA). During outlier exposure fine-tuning, Bilevel Prompt Distribution Learning (BPDL) is incorporated to optimize the pixel manifold distribution of the per-pixel mask embeddings. See Fig. 6 for details of BPDL. The image encoder extracts multi-scale panoramic image features in inference, which are then fed into the pixel decoder to generate semantically enriched features alongside the per-pixel embeddings for segmentation. Concurrently, textual prompts are embedded through the text encoder and projected into pixel space. These prompt embeddings, along with the decoded image features, are jointly processed by the transformer decoder, where they interact iteratively with learned query features. The refined queries are subsequently used for mask prediction and category classification. Finally, an OoD scoring mechanism is applied to produce the panoramic OoD segmentation prediction.
@@ III-B Framework
In contrast to pinhole-OoS, the complex backgrounds and pixel distortions in panoramic images hinder the generalization performance of pure visual models pre-trained on pinhole image datasets. This domain adaptation barrier prevents the model from effectively learning the visual features of panoramic images, resulting in suboptimal detection performance. It is important to explore methods for transferring models trained in the pinhole domain to the panoramic domain. Despite the strong performance of the mask-transformer model [29] in pinhole-OoS, the current advanced masked pinhole-OoS methods do not translate effectively to PanOoS. Moreover, accurate pixel decoding is essential for panoramic segmentation. During the critical outlier exposure fine-tuning, a conflict arises between the query embedding, which favors semantic classification, and the pixel decoding, which focuses on image details. Therefore, we optimize the architecture and fine-tuning procedures, introducing POS, the first mask-transformer-based segmentation model for PanOoS, as illustrated in Fig. 4. The main components of the model include four parts: the image encoder, the text encoder, the pixel decoder, and the Transformer decoder.
@@
Image Encoder. We utilize the image encoder from CLIP [31] for extracting image features and freeze its weights. The image encoder is configured as a ConvNext-Large [75] model, comprising four stages. It accepts an RGB image $x{\in}\mathbb{R}^{3\times H\times W}$ as input and outputs feature maps at several resolutions to the pixel decoder. Specifically, the output feature maps are downsampled with strides of $4(c_{4})$, $8(c_{8})$, $16(c_{16})$, and $32(c_{32})$ relative to the input image.
@@
Text Encoder. For the textual features, the semantic labels with $K$ classes and the distribution labels: inlier and outlier, are integrated into the pre-designed prompt templates, e.g.a photo of a {car}”, and “a photo of an {outlier} item in the urban or rural scenes”, to obtain the text descriptions about classes and distributions. The text encoder is structured as a $16$-layer transformer, each layer being $768$ units wide and featuring $12$ attention heads. The text descriptions are fed into text encoder to generate the prompt embeddings $\{T_{i}{\mid}i{=}0,\dots,K\}\bigcup\{P_{in},P_{out}\}{\in}\mathbb{R}^{(K+3)\times 768}$, while $T_{0}$ represents the learnable void embedding. Further, a linear projection is used to project the prompt embeddings into the pixel space to match the dimension of the mask feature. See the Appendix A-A for more details on prompt templates.
@@
Pixel Decoder. Following Mask2Former [29], the pixel decoder mainly consists of $6$ layers of Deformable Attention Transformer (DeformAttn) [76] and a Feature Pyramid Network (FPN) [77]. To facilitate out-of-distribution detection, the enriched semantic feature map with a stride of $32(c_{32}{})$ is processed with DeformAttn layers to produce $f_{4}$, which is then passed to the Transformer decoder. Then, pixel decoder gradually unsamples and fuses feature maps ($f_{4}$, $c_{16}$, $c_{8}$, and $c_{4}$) to generate per-pixel mask embeddings $F_{m}(x){\in}\mathbb{R}^{C_{m}\times H\times W}$, where $C_{m}$ is the embedding dimension.
@@
Transformer Decoder. The learnable query features $Q{\in}\mathbb{R}^{N\times C_{q}}$ are fed into the Transformer decoder layers, where they are jointly processed with the feature map $f_{4}$ and the prompt embeddings. $N$ represents the number of object queries and $C_{q}$ denotes the embedding dimension. Each decoder layer consists of a Prompt-based Restoration Attention followed by Self-attention, and a Feed-forward Network (FFN). Learnable positional embeddings are added to the query features.
@@
Mask and region class prediction. After refining the object queries and mask feature, the refined object queries are processed through a $3$-layer MLP to obtain $Q$ for predicting $N$ regions. Then, the membership score for each pixel belonging to a region is derived by multiplying $Q$ with the per-pixel mask feature $F_{m}(x)$ and applying a sigmoidsigmoid function:
$M(x)=\sigma(QF_{m}(x)).$ (2)
@@
In parallel, the refined object queries are fed into a linear layer followed by a softmaxsoftmax function to produce posterior class probabilities $P(x){\in}\mathbb{R}^{N\times K}$ for $K$ classes:
$P(x)=softmax(Q).$ (3)
@@
OoD Scoring. In contrast to pixel-wise semantic segmentation, the ground truth masks are divided into multiple binary masks, each containing all pixels belonging to a specific class. Ultimately, the class probabilities are broadcast to all pixels within the region, and the class score $S(x)$ is calculated as the product of the class probabilities and the mask predictions. Due to the unique architecture of the mask transformer [29, 71, 27], the class score $S(x)$ aggregates the weighted voting scores from all object queries; after training converges, specific object queries specialize in predicting a specific class. Therefore, the outlier scoring function from Rejected by All [27] is adopted:
$A(x)=RbA(S(x))=-\sum_{k=1}^{K}\phi(P(x)^{T}\!M(x)),$ (4)
@@
with $\phi$ being the tanhtanh activation function. $A(x)$ represents the outlier score map, indicating the outliers in the image.
@@
Panoramic-oriented Disentanglement. Existing research indicates that during fine-tuning, a balance must be maintained between mask segmentation and semantic prediction within the mask transformer [27]. PanOoS faces a similar conflict, and the decoding of panoramic per-pixel masks is crucial for the segmentation performance under conditions of semantic clarity. To disentangle the conflict between semantic query embedding and per-pixel mask decoding, we optimize the universal outlier exposure fine-tuning strategies, introducing a disentanglement strategy specifically designed for PanOoS. We observe that the performance of closed-set semantic segmentation for panoramic images during fine-tuning is crucial for identifying outliers. Specifically, the precise decoding of foreground pixels (i.e., masks) determines the model’s capability to detect outliers. Thus, to fully unleash the potential of the designed model for robust segmentation, we only fine-tune the pixel decoder, mask and region class prediction of POS during the outlier exposure phase.
@@ III-C Prompt-based Restoration Attention
In researching semantic learning for panoramic images, we found that traditional query-based decoding is difficult to effectively capture the intricate semantic information. This limitation primarily arises from prevalent pixel distortion and background clutter in images, which interfere with the semantic expression of queries, introduce semantic confusion, and ultimately impair the model’s comprehension of image content and segmentation performance. To address this challenge, we propose a Prompt-based Restoration Attention (PRA) module that utilizes the clear semantics embedded in text prompts to guide the semantic learning of queries. A semantic self-adaptive mechanism is proposed that dynamically adjusts the learning process based on the semantic information of queries, optimizing semantic extraction and generalization.
As shown in Fig. 5, we optimize the semantic learning process for queries. Specifically, we introduce an additional Cross Attention mechanism between the class and distribution prompt embeddings and the queries, built upon the existing Masked Attention. These prompt embeddings, with their clear semantic representation, provide robust semantic guidance for the query process. By incorporating the semantic information from the prompt embeddings, we mitigate potential semantic biases in the queries during the learning process, thereby significantly enhancing the semantic accuracy of queries. The interaction between the queries, image features, and prompt embeddings is formally expressed as:
$\hat{X}_{i,l}$ $=softmax(\mathcal{M}_{l-1}+Q_{l-1}~K^{T}_{i,l})~V_{i,l},$ (5)
$\hat{X}_{l}$ $=softmax(Q_{i}~K^{T}_{t})~V_{t},$ (6)
where $l$ is the layer index and $\hat{X}_{l}{\in}\mathbb{R}^{N\times C_{q}}$ refers to the enriched semantic queries at the $l^{th}$ layer. $Q_{l-1}$, $Q_{i}{\in}\mathbb{R}^{N\times C_{q}}$ are obtained by linearly transforming the previous queries $X_{l-1}$ and $\hat{X}_{i,l}{\in}\mathbb{R}^{N\times C_{q}}$ with learnable transformations, respectively. $\mathcal{M}_{l-1}$ denotes the attention mask [29]. $K_{i,l},V_{i,l}{\in}\mathbb{R}^{H_{l}W_{l}\times C_{q}}$ are the image features under learnable transformations respectively, and $H_{l},W_{l}$ are the spatial resolution of image features. $K_{t},V_{t}{\in}\mathbb{R}^{(K+3)\times C_{q}}$ are the prompt embeddings extracted by the text encoder under learnable transformations, respectively.
@@
Figure 5: Prompt-based Restoration Attention continuously employs masked attention and cross attention between image features and prompt embeddings while enhancing semantic clarity through query semantic self-adaptive correction.
@@
To further enhance the adaptability of query semantics, we propose a semantic self-adaptive correction mechanism, which dynamically adjusts the semantic representation of the query, allowing it to better adapt to complex pixel distributions and background information in the image, formally expressed as:
$X_{l}=\hat{X}_{l}+FFN(LN(\hat{X}_{l}))\cdot tanh(\mathcal{W}).$ (7)
@@
Here, LayerNorm (LN) is applied, and $\mathcal{W}$ represents the learnable adaptive weight. We initialize the weight value to zero to ensure the adaptation to diverse multi-scale image features while retaining the training stability. Through the adaptive mechanism, the query dynamically adjusts its representation based on learned semantic information.
@@ III-D Bilevel Prompt Distribution Learning
Figure 6: Bilevel Prompt Distribution Learning. The distribution prompts, text prompts for each class, and pixel embeddings are represented by stars, diamonds, and circles, respectively. The dashed circles indicate equidistant lines from the semantic center, and the dashed lines represent the decision boundary separating in- and out-of-distribution regions.
The pervasive pixel distortions and background clutter in panoramic images often result in a dispersed pixel manifold. This dispersion adversely affects the performance of deep neural networks, which tend to map the unknown samples inside the known class clusters instead of an arbitrary distribution in the whole feature space. Consequently, this misalignment distorts the voting scores of the pixel manifold, particularly in out-of-distribution regions within known classes, ultimately leading to overconfident and unreliable predictions. CLIP, which is trained on large-scale vision-language data pairs, demonstrates robust cross-domain generalization capabilities. Moreover, its pretrained text encoder adeptly constructs the semantic feature distribution of known classes while maintaining consistency with the pixel space distribution. This accurate semantic feature distribution provides a reliable reference for pixel manifold learning. Leveraging this foundation, we propose a Bilevel Prompt Distribution Learning (BPDL) loss, which utilizes projected class and distribution prompts $\{T_{i}{\mid}i{=}0,\dots,K\}\bigcup\{P_{in},P_{out}\}{\in}\mathbb{R}^{(K+3)\times C_{m}}$, to optimize the spatial distribution of per-pixel embeddings $F_{m}(x)$.
As shown in Fig. 6, its core objective is to enhance the discriminability of out-of-distribution embeddings $\{O_{i}{\mid}i{=}1,\dots,N_{o}\}{\in}\mathbb{R}^{N_{o}\times C_{m}}$ at the decision boundary, thereby improving the model’s capability to accurately identify out-of-distribution regions. Specifically, the bilevel prompt distribution learning optimizes per-pixel embeddings at two levels (in- and out-of-distribution) across four key aspects.
In-distribution learning. To prevent the model from misinterpreting intra-class pixel embeddings when distinguishing out-of-distribution from in-distribution pixel embeddings $\{m_{i}{\mid}i{=}1,\dots,N_{i}\}{\in}\mathbb{R}^{N_{i}\times C_{m}}$, thereby improving closed-set segmentation performance, we introduce an in-distribution loss $\mathcal{L}_{pixel}$. As shown in Fig. 7, it guides the model’s learning process by enhancing intra-class pixel embedding compactness while preserving inter-class separability. First, we minimize the Euclidean distance between each pixel embedding and the corresponding class center, optimizing intra-class compactness, which can be represented as follows:
$\mathcal{L}_{intra}=\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}\left\|m_{i}-T_{k_{i}}\right\|_{2}^{2},$ (8)
where $k_{i}$ represents the class label of the in-distribution pixel embedding $m_{i}$.

### v3

@@ III Methodology
In this work, we introduce a novel panoramic perception task, PanOoS, to promote a more comprehensive omnidirectional perception by leveraging the broader FoV of panoramic imagery. In response to the challenges posed by pixel distortions and complex backgrounds, we analyze the model architecture and fine-tuning strategies, ultimately proposing the first mask-based out-of-distribution segmentation model tailored to the PanOoS task.
@@ III-A Problem Setting
Out-of-distribution segmentation generalizes anomaly segmentation by requiring not only the detection of anomalous pixels but also accurate semantic labeling of all in-distribution categories. Formally, given an input image $\mathbf{x}{\in}\mathbb{R}^{3\times H\times W}$, the pixel-wise segmentation architectures [cheng2022masked, nayal2023rba, rai2024mask2anomaly, grcic2023advantages, delic2024outlier] produce per-pixel class logits $\mathbf{S}(x){\in}\mathbb{R}^{K\times H\times W}$, where $K$ denotes the number of classes for closed-set segmentation, and $H$ and $W$ correspond to the height and width of the image, respectively. An anomaly score map $\mathbf{A}(x)$ is defined as
$\mathbf{A}(x)=f(\mathbf{S}(x))\in\mathbb{R}^{H\times W},$ (1)
where $f{:}~\mathbb{R}^{K}\to\mathbb{R}$ maps the per-pixel class logits to a scalar out-of-distribution confidence score.
Figure 4: Overview of POS. The proposed POS consists of four key components: an image encoder, a text encoder, a pixel decoder, and a transformer decoder with Prompt-based Restoration Attention (PRA). During outlier exposure fine-tuning, we incorporate Bilevel Prompt Distribution Learning (BPDL) to optimize the pixel manifold distribution of the per-pixel mask embeddings. See Fig. 6 for details of BPDL. During inference, the image encoder extracts multi-scale image features, which are fed into the pixel decoder to generate semantically enriched features alongside the per-pixel embeddings for segmentation. Concurrently, textual prompts are embedded through the text encoder and projected into pixel space. These prompt embeddings, along with the decoded image features, are jointly processed by the transformer decoder, where they interact iteratively with learned query features. The refined queries are subsequently used for mask prediction and category classification. Finally, an OoD scoring mechanism is applied to produce the out-of-distribution segmentation prediction.
@@ III-B Framework
Despite the strong performance of the mask-transformer model [cheng2022masked] in pinhole-OoS, the current methods fail to generalize to PanOoS. The domain adaptation barrier prevents the model from effectively learning the visual features of panoramic images, resulting in suboptimal segmentation performance. It is important to explore methods for transferring models trained in the pinhole domain to the panoramic domain. Moreover, accurate pixel decoding is essential for panoramic segmentation. During the critical outlier exposure fine-tuning, a conflict arises between the query embedding, which favors semantic classification, and the pixel decoding, which focuses on image details. Therefore, we optimize the architecture and fine-tuning procedures, introducing POS, the first mask-transformer-based segmentation model for PanOoS, as illustrated in Fig. 4. The model comprises four main components: the image encoder, the text encoder, the pixel decoder, and the Transformer decoder.
III-B1 Image Encoder
@@
We utilize the image encoder from CLIP [radford2021learning] for extracting image features and freeze its weights. The image encoder is a ConvNeXt-Large model pretrained in the CLIP framework [liu2022convnet], comprising four stages. It accepts an RGB image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$ as input and outputs feature maps at several resolutions to the pixel decoder. Specifically, the output feature maps are downsampled with strides of $4(\mathbf{c}_{4})$, $8(\mathbf{c}_{8})$, $16(\mathbf{c}_{16})$, and $32(\mathbf{c}_{32})$ relative to the input image.
III-B2 Text Encoder
@@
For the textual features, the $K$ semantic class names and distribution nouns, inlier and outlier, are integrated into the pre-designed prompt templates to obtain text descriptions about classes and distributions. The text encoder is structured as a $16$-layer transformer, each layer being $768$ units wide and featuring $12$ attention heads. The descriptions are fed into the text encoder to generate the prompt embeddings $\{\mathbf{T}_{i}\mid i=0,\dots,K\}\cup\{\mathbf{P}_{in},\mathbf{P}_{out}\}{\in}\mathbb{R}^{(K+3)\times 768}$, where $\mathbf{T}_{0}$ denotes the learnable void embedding. A linear projection is then applied to project the prompt embeddings into the pixel space to match the dimensionality of per-pixel embeddings. Prompt template details are in the Appendix.
III-B3 Pixel Decoder
@@
Following Mask2Former [cheng2022masked], the pixel decoder comprises $6$ layers of Deformable Attention Transformer (DAT) [zhu2020deformable] and a Feature Pyramid Network (FPN) [lin2017feature]. To facilitate OoS, the enriched semantic feature map $\mathbf{c}_{32}$ is processed by the DAT layers to produce $\mathbf{f}_{4}$, which is then passed to the transformer decoder. Then, the pixel decoder progressively upsamples and fuses feature maps ($\mathbf{f}_{4}$, $\mathbf{c}_{16}$, $\mathbf{c}_{8}$, and $\mathbf{c}_{4}$) to generate per-pixel mask embeddings $\mathbf{F}_{m}{\in}\mathbb{R}^{C_{m}\times H\times W}$, where $C_{m}$ is the embedding dimension.
III-B4 Transformer Decoder
@@
The learnable query features $\mathbf{Q}{\in}\mathbb{R}^{N\times C_{q}}$ are fed into the Transformer decoder, where they are jointly processed with $\mathbf{f}_{4}$ and the prompt embeddings. $N$ and $C_{q}$ denote the number of object queries and the embedding dimension, respectively. Each decoder layer consists of a Prompt-based Restoration Attention, followed by self-attention and a Feed-Forward Network (FFN). Learnable positional embeddings are added to the query features.
III-B5 Mask and region class prediction
@@
After refining the per-pixel embeddings and object queries, $\mathbf{Q}$ is passed through a $3$-layer MLP to obtain per-segment embeddings for mask prediction. The membership score for each pixel belonging to a region is derived by multiplying the per-segment embeddings with per-pixel embeddings $\mathbf{F}_{m}$ and applying a sigmoid function:
$\mathbf{M}(x)=\sigma(\mathrm{MLP}(\mathbf{Q})~\mathbf{F}_{m})\in\mathbb{R}^{N\times H\times W}.$ (2)
@@
In parallel, the refined object queries are fed into a linear layer followed by a softmax function to produce posterior class probabilities $\mathbf{P}(x)$ of $K$ classes:
$\mathbf{P}(x)=\mathrm{softmax}(\mathbf{Q}\mathbf{W}_{p}^{\top})\in[0,1]^{N\times K},$ (3)
where $\mathbf{W}_{p}{\in}\mathbb{R}^{K\times C_{q}}$ is a learnable weight matrix.
III-B6 OoD Scoring
@@
In contrast to per-pixel classification, the ground truth masks are divided into multiple binary masks, each containing all pixels belonging to a specific class. Ultimately, the class probabilities $\mathbf{P}(x)$ are broadcast to all pixels within the region, and the class logits $\mathbf{S}(x)$ are calculated as the product of the mask predictions $\mathbf{M}(x)$ with $\mathbf{P}(x)$:
$\mathbf{S}(x)=\sum_{n=1}^{N}\mathbf{P}_{n}(x)\mathbf{M}_{n}(x).$ (4)
Due to the unique architecture of mask transformer [cheng2022masked, ackermann2023maskomaly, nayal2023rba], the class logits $\mathbf{S}(x)$ aggregate weighted voting scores from all object queries; upon convergence, object queries specialize in predicting a specific class. Thus, the outlier scoring function from Rejected by All [nayal2023rba] is adopted:
$\mathbf{A}(x)=\mathrm{RbA}(\mathbf{S}(x))=-\sum_{k=1}^{K}\phi(\mathbf{S}_{k}(x)),$ (5)
@@
with $\phi$ being the tanhtanh activation function. $\mathbf{A}(x)$ represents the outlier score map, indicating the outliers in the image $\mathbf{x}$.
III-B7 Panoramic-oriented Disentanglement
@@
Existing research indicates that during fine-tuning, a balance must be maintained between mask segmentation and semantic prediction within the mask transformer [nayal2023rba]. We observe that PanOoS faces a similar conflict, and accurate decoding of panoramic per-pixel masks is crucial for segmentation performance under conditions of semantic clarity. Specifically, the precise decoding of foreground pixels (i.e., masks) determines the model’s capability to detect outliers. To disentangle the conflict between semantic query and per-pixel mask decoding, we introduce a disentanglement strategy specifically designed for PanOoS, and only fine-tune the pixel decoder, the $3$-layer MLP, and the linear layers of POS during the outlier exposure phase.
@@ III-C Prompt-based Restoration Attention
Traditional query-based decoding is ineffective at capturing the intricate semantic information of panoramic imagery. This limitation primarily arises from prevalent pixel distortion and background clutter in images, which interfere with the semantic expression of queries, introduce semantic confusion, and ultimately impair the model’s comprehension of image content and segmentation performance. To address this challenge, we propose a Prompt-based Restoration Attention (PRA) module that utilizes the clear semantics embedded in text prompts to guide the semantic learning of queries. A self-adaptive mechanism that dynamically adjusts the learning process based on the semantic information of queries is proposed to enhance the robustness and generalization of semantic extraction.
@@
Figure 5: Prompt-based Restoration Attention continuously employs masked attention and cross attention between image features and prompt embeddings while refining semantic clarity through query semantic self-adaptive correction.
Specifically, as shown in Fig. 5, we introduce an additional Cross Attention mechanism between both class and distribution prompt embeddings and the queries, built upon the existing Masked Attention. These prompt embeddings, with their clear semantic representations, provide robust semantic guidance for the query process. By incorporating the semantic information from the prompt embeddings, we mitigate potential semantic biases in the queries during the learning process, thereby significantly enhancing the semantic fidelity of queries. The interaction between the queries, image features, and prompt embeddings is formally expressed as:
$\hat{\mathbf{X}}_{i,l}$ $=\mathrm{softmax}(\mathcal{M}_{l-1}+\mathbf{Q}_{l-1}\mathbf{K}_{i,l}^{T})\mathbf{V}_{i,l},$ (6)
$\hat{\mathbf{X}}_{l}$ $=\mathrm{softmax}(\mathbf{Q}_{i}\mathbf{K}_{t}^{T})\mathbf{V}_{t},$ (7)
where $l$ denotes the layer index, and $\hat{\mathbf{X}}_{l}{\in}\mathbb{R}^{N\times C_{q}}$ represents the enriched semantic queries at the $l^{th}$ layer. $\mathbf{Q}_{l-1}$ and $\mathbf{Q}_{i}{\in}\mathbb{R}^{N\times C_{q}}$ are obtained by applying learnable linear transformations to the previous queries $\mathbf{X}_{l-1}$ and $\hat{\mathbf{X}}_{i,l}{\in}\mathbb{R}^{N\times C_{q}}$, respectively. $\mathcal{M}_{l-1}$ denotes the attention mask [cheng2022masked]. $\mathbf{K}_{i,l},\mathbf{V}_{i,l}{\in}\mathbb{R}^{H_{l}W_{l}\times C_{q}}$ are the image features under learnable transformations respectively, and $H_{l},W_{l}$ are the spatial resolution of image features. $\mathbf{K}_{t},\mathbf{V}_{t}{\in}\mathbb{R}^{(K+3)\times C_{q}}$ are the prompt embeddings extracted by the text encoder under learnable transformations, respectively.
@@
Then, we propose a semantic self-adaptive correction mechanism, which dynamically refines the semantic representation of the query, formally expressed as:
$\mathbf{X}_{l}=\hat{\mathbf{X}}_{l}+\mathrm{FFN}(\mathrm{LN}(\hat{\mathbf{X}}_{l}))\cdot\tanh(\mathcal{W}).$ (8)
@@
Here, LayerNorm (LN) is applied to $\hat{\mathbf{X}}_{l}$, and $\mathcal{W}$ represents the learnable adaptive weight. We initialize the weight value to zero to ensure the adaptation to diverse multi-scale image features while retaining training stability.
@@ III-D Bilevel Prompt Distribution Learning
Figure 6: Bilevel Prompt Distribution Learning. The distribution prompts, text prompts for each class, and pixel embeddings are denoted by stars, diamonds, and circles, respectively. The dashed circles indicate equidistant lines from the semantic center, and the dashed lines represent the decision boundary separating in- and out-of-distribution regions.
The pervasive pixel distortions and background clutter in panoramic images often result in dispersed pixel manifolds. This dispersion adversely affects the performance of deep neural networks, which tend to map the unknown samples inside the known class clusters instead of an arbitrary distribution in the whole feature space. Consequently, this misalignment distorts the voting scores of pixel manifolds, particularly in out-of-distribution regions, ultimately leading to overconfident and unreliable predictions. CLIP, trained on large-scale vision-language data pairs, demonstrates robust cross-domain generalizability. Its pretrained text encoder adeptly captures the semantic feature distribution of known classes while maintaining consistency with the pixel space distribution. This accurate semantic feature distribution provides a reliable reference for pixel manifold learning. Leveraging this foundation, we propose a Bilevel Prompt Distribution Learning (BPDL) loss that utilizes projected class and distribution prompts $\{\mathbf{T}_{i}\mid i=0,\dots,K\}\cup\{\mathbf{P}_{in},\mathbf{P}_{out}\}\in\mathbb{R}^{(K+3)\times C_{m}}$ to optimize the spatial distribution of per-pixel embeddings $\mathbf{F}_{m}$.
As shown in Fig. 6, its core objective is to enhance the discriminability of out-of-distribution pixel embeddings $\{\mathbf{O}_{i}\mid i=1,\dots,N_{o}\}\in\mathbb{R}^{N_{o}\times C_{m}}$ near the decision boundary. Specifically, the BPDL optimizes per-pixel embeddings at two levels (in- and out-of-distribution) across four key aspects.
III-D1 In-distribution learning
To improve closed-set segmentation performance, particularly by maintaining coherent intra-class representations when distinguishing out-of-distribution from in-distribution pixel embeddings $\{\mathbf{m}_{i}\mid i=1,\dots,N_{i}\}\in\mathbb{R}^{N_{i}\times C_{m}}$, we introduce an in-distribution loss $\mathcal{L}_{pixel}$. First, as shown in Fig. 7, we minimize the Euclidean distance between each pixel embedding and its corresponding class center to enhance the intra-class compactness:
$\mathcal{L}_{intra}=\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}\left\|\mathbf{m}_{i}-\mathbf{T}_{k_{i}}\right\|_{2}^{2},$ (9)
where $k_{i}$ denotes the class label of the embedding $\mathbf{m}_{i}$.